# TCKIN: A Novel Integrated Network Model for Predicting Mortality Risk in Sepsis Patients

### v1

@@ Abstract
Sepsis poses a major global health threat, accounting for millions of deaths annually and significant economic costs. Accurate predictions of mortality risk in sepsis patients facilitate the efficient allocation of medical resources, thereby enhancing patient survival and quality of life. Through precise risk assessments, healthcare facilities can effectively distribute intensive care beds, medical equipment, and staff, ensuring high-risk patients receive timely and appropriate care. Early identification and intervention significantly decrease mortality rates and improve patient outcomes. Current methods typically utilize only one type of data—either constant, temporal, or ICD codes. This study introduces the Time-Constant KAN Integrated Network (TCKIN), an innovative model that enhances the accuracy of sepsis mortality risk predictions by integrating both temporal and constant data from electronic health records and ICD codes. Validated against the MIMIC-III and MIMIC-IV datasets, TCKIN surpasses existing machine learning and deep learning methods in accuracy, sensitivity, and specificity. Notably, TCKIN achieved AUCs of 87.76% and 88.07%, demonstrating superior capability in identifying high-risk patients. Additionally, TCKIN effectively combats the prevalent issue of data imbalance in clinical settings, improving the detection of patients at elevated risk of mortality and facilitating timely interventions. These results confirm the model’s effectiveness and its potential to transform patient management and treatment optimization in clinical practice. With this advanced risk assessment tool, healthcare providers can devise more tailored treatment plans, optimize resource utilization, and ultimately enhance survival rates and quality of life for sepsis patients.
@@
Sepsis , Mortality Risk Prediction , GRU-D , KAN
@@
organization=Yunnan University,addressline=dongfanglin@stu.ynu.edu.cn, city=Kunming, postcode=650000, state=Yunnan Province, country=China
@@ 1 Introduction
Sepsis, a systemic inflammatory response syndrome triggered by infection [1], can lead to multi-organ dysfunction or failure in severe cases [2]. As a significant global health challenge, sepsis affects millions annually, resulting in substantial medical costs and high mortality rates [3]. According to the World Health Organization, approximately 49 million cases occur worldwide each year, with about 11 million fatalities [4]. The economic impact is profound; for instance, in the United States, the annual cost of treating sepsis exceeds \$24 billion, representing one of the healthcare system’s largest expenditures [5]. Given the high mortality associated with sepsis [6] and its economic repercussions, accurately predicting mortality risk in sepsis patients is crucial. This enables more rational allocation of medical resources, enhancing patient survival rates and quality of life [7]. Through precise mortality risk assessments, medical institutions can more effectively distribute resources like intensive care beds and medical staff, ensuring high-risk patients receive timely care. Accurate predictions also allow for the development of personalized treatment plans, optimizing treatment strategies to improve efficacy, reduce hospital stays, and lower costs [8]. Most critically, early identification and intervention for high-risk patients can significantly decrease mortality rates, thereby improving outcomes and quality of life.
With the advent of Electronic Health Records (EHR) systems [9], there has been a significant increase in the availability of medical data, enriching the resources for mortality risk prediction in sepsis. Data typically used for this purpose includes physiological metrics from the prior 24 hours, laboratory results, and demographic and diagnostic information, which are pivotal for developing high-accuracy predictive models.
In current clinical practice, several scoring tools are widely utilized to assess sepsis severity, including the Sequential Organ Failure Assessment (SOFA) [10], Acute Physiology and Chronic Health Evaluation II (APACHE II) [11], Simplified Acute Physiology Score II (SAPS II) [12], and the Sepsis Severity Score (SSS) [13]. However, these tools exhibit notable limitations. They are predominantly based on specific populations and clinical settings from their time of development, leading to reduced predictive accuracy as demographics and practices evolve. Studies indicate that SOFA, APACHE II, and SAPS II are less effective in predicting ICU sepsis patient mortality due to their reliance on a limited set of static physiological parameters, which do not fully capture dynamic patient condition changes [14, 15]. Furthermore, although the Sepsis Severity Score is tailored for in-hospital sepsis severity, it similarly shows weak predictive performance for in-hospital mortality [16].
Advancements in machine learning have spurred the use of its algorithms to enhance sepsis mortality predictions. For example, Taylor et al. employed a local big data-driven Random Forest model [17], using over 500 clinical variables to predict emergency department sepsis patient mortality, yielding results superior to those of Logistic Regression [18]. Continuing this trend, Zhang et al. developed a LASSO score using the MIMIC-III dataset [19], although it still requires external validation. Additionally, Kong et al. created four machine learning models, including LASSO [20], Random Forest [17], Gradient Boosting Machine [21], and Logistic Regression, using 86 features from the MIMIC-III dataset to predict in-hospital mortality in ICU sepsis patients. These models, except for Random Forest—which underestimated high-risk patient mortality—showed good calibration, in contrast to the poor calibration of the SAPS II score [22]. However, traditional machine learning techniques face challenges with handling extensive feature sets and time-series data, potentially leading to overfitting and limited effectiveness in sepsis mortality prediction [23, 24].
@@
Deep learning, a pivotal branch of machine learning, has significantly advanced the medical field recently. Its applications span medical image analysis [25], genetic data analysis [26], and disease risk prediction[27]. Deep learning surpasses traditional methods with its robust feature extraction and pattern recognition capabilities, enabling the automatic learning of useful features from extensive, complex datasets. Additionally, deep learning effectively processes time-series data, thus enhancing its predictive accuracy in such environments. In a study on 30-day mortality rates for sepsis patients, Su et al. employed a basic two-layer multilayer perceptron (MLP) model using only 11 static data features[28]. Although this method achieved some predictive success, it failed to incorporate patients’ temporal data, which restricted its ability to detect time-dependent changes. Moreover, the model’s simplicity likely limited its capacity to uncover complex data patterns, thereby capping further improvements in predictive performance. In subsequent research, Cheng et al. utilized three machine learning models—Convolutional Neural Network (CNN)[29], Long Short-Term Memory (LSTM) [30], and Random Forest (RF)—to analyze sepsis patients’ temporal vital signs data, predicting in-hospital mortality[31]. However, their study was confined to a minimal set of seven features, which might have overlooked more complex variables influencing patient outcomes. Similarly, Gong et al. developed a Temporal Convolutional Networks (TCN)-based model that integrated time-series data with four vital signs to predict near-term mortality risks in sepsis patients, yet it also limited its analysis to a few variables[32]. Remarkably, none of these studies considered the impact of patients’ diagnostic ICD codes on mortality risk predictions, which could be a critical oversight.
In this study, we introduce the Temporal-Constant KAN Integrated Network (TCKIN) designed to predict the mortality risk of ICU sepsis patients. Figure 1 illustrates the main flowchart of the TCKIN model. Our principal contributions are outlined as follows:
Figure 1: main flowchart of the TCKIN
@@
    Integration of Diagnostic ICD Codes and CCS Medical Ontology: We have integrated the patient’s diagnostic ICD codes with the Clinical Classifications Software (CCS) medical ontology into our model using graph networks. This integration significantly enhances the model’s predictive accuracy by fully leveraging these coding systems.
@@
    Fusion of Temporal and Constant Data: Our approach merges temporal data handling via the GRU-D[33] with constant data analysis performed by the KAN network[34]. The KAN network extracts critical information from constant data, which is then combined with the temporal data processed by GRU-D. This combined data, along with ICD and CCS codes, undergoes further processing through the KAN network to notably improve the model’s predictive capabilities.
@@
    Superior Model Performance: Experimental results demonstrate that the TCKIN substantially surpasses existing methods in prediction accuracy and robustness. By integrating multiple data sources and processing techniques, TCKIN achieves exceptional performance across various evaluation metrics, thereby offering more reliable support for clinical decision-making.
@@ 2.2 Data Preprocessing
Based on the Sepsis-3 definition[36], we selected sepsis patients from the MIMIC-III and MIMIC-IV datasets who had a SOFA score of 2 or higher, as illustrated in Figure 2. Initially, we excluded patients under the age of 16 and those with corrupted data. We also eliminated patients with suspected infections occurring more than 24 hours before or after ICU admission, due to the ambiguous timing of their infection diagnoses, which could compromise the study’s accuracy. Furthermore, we excluded patients with ICU stays shorter than 24 hours, as their incomplete clinical data insufficiently support the model. Additionally, patients who underwent cardiac, thoracic, or vascular surgery were omitted because their unique surgical treatments and care requirements could adversely affect the model’s generalizability. For sepsis patients with multiple hospital admissions, we used data from their initial admission.
@@
Using these criteria, we extracted patients’ constant and temporal data, along with diagnostic ICD coding information, from the MIMIC-III and MIMIC-IV databases. Constant data includes basic demographic information like gender, age, and weight, which provides essential background on the patients. Temporal data encompasses detailed physiological signs and laboratory test results recorded during the ICU stay, such as temperature, heart rate, blood pressure, respiratory rate, hemoglobin levels, and white blood cell count, reflecting the immediate health status and condition changes of the patients. Additionally, diagnostic ICD codes were transformed into vector representations via graph embedding methods, allowing for a comprehensive integration of patients’ constant characteristics, temporal features, and diagnostic information. Table 1 provides a detailed listing of the features extracted from the temporal and constant data.
@@
Temporal Feature Platelet, MCH, MCHC, Fibrinogen, Basophils, Lymphocytes, HCO3, Troponin-t, Sodium, Urine output, Bilirubin, HR, Lactic acid, Albumin, Calcium, Glucose, PH, Creatinine, Eosinophils, Total CO2, SBP, RBC, Chloride, ALT, Fraction inspired oxygen set, Alkaline phosphate, PaO2, GCS, PT, SpO2, Anion gap, DBP, Magnesium, BUN, WBC, Glucose, Temperature, MCV, Troponin-i, Potassium, Neutrophils, Hemoglobin, PTT, Lactate, PaCO2, FiO2, Hematocrit, AST, Monocytes, Weight, Phosphate, INR, MBP, RR
@@
In preprocessing the constant data extracted from patients, we implemented specific strategies for different types of data missingness. For continuous variables such as age and weight, missing values were filled with zero to maintain data completeness and consistency without introducing bias, ensuring that the model’s training is unaffected by missing values. For categorical variables such as gender and race, we employed one-hot encoding, which converts these variables into binary vectors where each category corresponds to a unique dimension. This approach prevents the imposition of implicit ordinal relationships among categories, thereby avoiding misleading influences on the model’s training. In processing the temporal data, we adhered to strict time window and interval strategies: temporal data were extracted within the 24 hours prior to ICU admission and sampled at 1-hour intervals. For each interval, single measurement values were used directly, while multiple measurements were averaged to represent the feature’s level during that hour. In cases where no measurements were recorded within an interval, values were designated as missing. Figure 3 illustrates the processing of temporal data. We will address the treatment of missing values in temporal data during subsequent model processing.
@@
$H^{(l+1)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$ (1)
@@ 3 Label
2.3 The Architecture of TCKIN
In our study, the proposed network architecture comprises three primary components. Initially, temporal data are processed via the GRU-D (Gated Recurrent Unit with Decay) model, which generates hidden representations of these data. The second component utilizes an attention mechanism to analyze the patient’s ICD diagnostic codes and CCS codes, capturing complex relationships and semantic information, thereby enriching diagnostic feature representations. The third component processes constant data through the Kolmogorov–Arnold Networks (KAN). This network efficiently leverages constant feature information to extract high-level features via a multi-layer neural network. Subsequently, the hidden features from these three components are concatenated to form an integrated feature vector. This vector is further processed through a final KAN network to predict the patient’s sepsis mortality risk.
@@ 2.3 The Architecture of TCKIN
$D$ $=(D_{1},D_{2},\dots,D_{t},\dots,D_{T})\in\mathbb{R}^{T\times N}$ (2)
@@ 2.4 Implementation
In our research, we implemented and trained our model using the TensorFlow framework. We utilized the Adam optimizer to update model parameters and employed a learning rate decay strategy to dynamically adjust the learning rate, enhancing the model’s convergence stability. To address the significant imbalance between positive and negative samples, we applied oversampling techniques to maintain a consistent ratio of these samples across training, validation, and test sets. This method effectively mitigated model bias toward the majority class, thereby enhancing its generalization capabilities on imbalanced datasets. Moreover, we adopted five-fold cross-validation to validate the model’s stability and generalizability. By dividing the dataset into five mutually exclusive subsets, using one for validation and the remaining four for training—and repeating this process five times—we ensured the model’s consistent performance across various data splits. To prevent overfitting, we incorporated Dropout layers and monitored the validation set performance using early stopping techniques, halting training when no further reduction in validation loss was observed.
@@ 3.1 Performance Metrics
In this section, we describe the metrics employed to assess the TCKIN model’s performance. These metrics are vital for gauging the model’s accuracy in predicting mortality risk among ICU sepsis patients. The principal evaluation metrics include sensitivity, specificity, area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), and the Brier Score (BS). Each metric offers a unique perspective on the model’s performance, enabling a thorough analysis. The specific formulas for these metrics are as follows:
@@ 3.2 Model Performance
We evaluated the TCKIN model’s performance on the MIMIC-III and MIMIC-IV datasets by comparing it with seven established baseline models: Xgboost[38], SVM, Random Forest (RF)[39], LGBM[22], LSTM[31], IseeU[40], and and Che et al.’s model GRU-D[33]. This comparison across various metrics such as sensitivity, specificity, AUROC, AUPRC, and Brier score allowed us to thoroughly assess the strengths and weaknesses of each model. Figure 6 illustrates the comparative performance, demonstrating that the TCKIN model generally outperforms the baseline models. Notably, on the MIMIC-IV dataset, the TCKIN model achieved AUROC and AUPRC values of 0.8807 and 0.5470, respectively, significantly surpassing its competitors. This indicates its superior accuracy and stability in distinguishing between positive and negative cases and handling imbalanced datasets. Additionally, the TCKIN model’s balanced performance in sensitivity (0.7875) and specificity (0.8153) underscores its effectiveness in identifying patients at different levels of risk. A lower Brier score of 0.0696 reflects the model’s precision in probability predictions. Figure 7, showing the ROC curves, further highlights the TCKIN model’s enhanced ability to differentiate between outcomes, with its curve closer to the top-left corner, symbolizing higher prediction accuracy across various thresholds.
@@
Figure 7: Receiver Operating Characteristic (ROC) curves, employing 5-fold cross-validation on MIMIC-III and MIMIC-IV datasets, display the performance of various models. These include GRUD, ISEEU, LightGBM, Logistic Regression (LR), Long Short-Term Memory (LSTM), Random Forest (RF), Support Vector Machine (SVM), TCKIN, and XGBoost.
To further evaluate the robustness of the TCKIN model, we conducted parameter sensitivity experiments focusing on the learning rate (LR) and batch size, two crucial hyperparameters that influence model performance. We systematically varied these parameters, setting the learning rate at 0.0001, 0.001, 0.01, and 0.1, and the batch size at 16, 32, 64, and 128, to observe their effects on model performance under various combinations. Figure 8 illustrates these performance changes through a three-dimensional graph that plots AUROC values against variations in learning rate and batch size. The results indicate that the TCKIN model achieves optimal performance with a learning rate of approximately 0.001 and a batch size of 64. Notably, excessively high or low learning rates can degrade performance, while very small batch sizes can introduce instability during training, and overly large batch sizes may reduce training speed and impair generalization.
Figure 8: The impact of varying batch sizes and learning rates on the performance of the TCKIN model across MIMIC-III and MIMIC-IV datasets. Each bar represents the Area Under the Receiver Operating Characteristic Curve (AUROC), with colors indicating different AUROC values as batch size and learning rate change.

### v2

# TCKAN: A Novel Integrated Network Model for Predicting Mortality Risk in Sepsis Patients

@@ Abstract
Sepsis poses a major global health threat, accounting for millions of deaths annually and significant economic costs. Accurately predicting the risk of mortality in sepsis patients enables early identification, promotes the efficient allocation of medical resources, and facilitates timely interventions, thereby improving patient outcomes. Current methods typically utilize only one type of data—either constant, temporal, or ICD codes. This study introduces a novel approach, the Time-Constant Kolmogorov-Arnold Network (TCKAN), which uniquely integrates temporal data, constant data, and ICD codes within a single predictive model. Unlike existing methods that typically rely on one type of data, TCKAN leverages a multi-modal data integration strategy, resulting in superior predictive accuracy and robustness in identifying high-risk sepsis patients. Validated against the MIMIC-III and MIMIC-IV datasets, TCKAN surpasses existing machine learning and deep learning methods in accuracy, sensitivity, and specificity. Notably, TCKAN achieved AUCs of 87.76% and 88.07%, demonstrating superior capability in identifying high-risk patients. Additionally, TCKAN effectively combats the prevalent issue of data imbalance in clinical settings, improving the detection of patients at elevated risk of mortality and facilitating timely interventions. These results confirm the model’s effectiveness and its potential to transform patient management and treatment optimization in clinical practice. Although the TCKAN model has already incorporated temporal, constant, and ICD code data, future research could include more diverse medical data types, such as imaging and laboratory test results, to achieve a more comprehensive data integration and further improve predictive accuracy.
@@
Sepsis Patient , Mortality Risk Prediction, Kolmogorov-Arnold Network
@@
[…]
@@ 1 Introduction
Sepsis, a life-threatening condition characterized by a systemic inflammatory response to infection[1], poses a significant global health burden[2, 3]. Accurate prediction of sepsis mortality risk is essential for efficient resource allocation and improved patient outcomes [4, 5, 6].
To effectively allocate resources and improve patient outcomes, numerous models have been developed to predict the mortality risk associated with sepsis. Scoring tools [7, 8] are widely used to assess the severity of sepsis. While various scoring tools have been developed, these tools are predominantly based on specific populations and clinical settings from their time of development, leading to reduced predictive accuracy as demographics and practices evolve. Additionally, these methods rely on a limited set of static physiological parameters that do not fully capture the dynamic changes in patient conditions[9, 10].
The rise of Electronic Health Records (EHRs) has enhanced the availability of medical data, crucial for developing predictive models using physiological, laboratory, and demographic information [11]. This has spurred the application of machine learning in sepsis mortality prediction. For instance, Taylor et al. demonstrated that a Random Forest model, utilizing over 500 clinical variables, outperformed Logistic Regression in predicting sepsis mortality [12, 13]. Similarly, Zhang et al. developed a LASSO-based score using the MIMIC-III dataset [14], though external validation is still required. Kong et al. extended this work by developing four models—LASSO [15], Random Forest [13], Gradient Boosting Machine [16], and Logistic Regression—based on 86 features from MIMIC-III to predict in-hospital mortality [17]. However, traditional machine learning methods often encounter challenges when dealing with large feature sets and time-series data, which can lead to overfitting and reduced effectiveness in sepsis mortality prediction.
@@
Deep learning, a pivotal branch of machine learning, has recently made significant strides in biomedicine, with applications in genomics [18, 19], metabolomics [20], and proteomics [21]. In health informatics, particularly in mortality prediction, deep learning has demonstrated its capability to automatically extract features from large datasets and handle time-series data, thereby enhancing predictive accuracy [22]. For instance, For example, Su et al. used a two-layer MLP model to predict 30-day mortality in sepsis patients, but the lack of temporal data limited its predictive power [23]. Cheng et al. utilized Convolutional Neural Network (CNN)[24], Long Short-Term Memory (LSTM) [25], and Random Forest (RF) to analyze sepsis patients’ temporal vital signs data, predicting in-hospital mortality[26]. However, their study was confined to a minimal set of seven features, which might have overlooked more complex variables influencing patient outcomes. Similarly, Gong et al. developed a Temporal Convolutional Networks (TCN)-based model that integrated time-series data with four vital signs to predict near-term mortality risks in sepsis patients, yet it also limited its analysis to a few variables[27]. Moreover, Recent studies support the benefits of integrating multiple data sources to improve predictive model performance [28]. and knowledge-guided approaches can improve diagnostic predictions [29]. However, existing studies often rely on limited sets of features, fail to effectively integrate temporal and constant data, or overlook the incorporation of domain-specific additional knowledge.
To tackle the above issues, this study introduces the Temporal-Constant Kolmogorov-Arnold Network (TCKAN) to enhance sepsis mortality prediction. Building upon the GRU-D[30] and KAN architectures[31], TCKAN effectively integrates temporal data, constant data, and diagnostic ICD codes into a unified model. Figure 1 illustrates the main flowchart of the TCKAN model. Our principal contributions are outlined as follows:
Figure 1: Main flowchart of the TCKAN
@@
    Employing graph neural networks to integrate diagnostic ICD codes and Clinical Classification Software (CCS) for a more informative representation of patient characteristics.
@@
    Developing an effective fusion mechanism using GRU-D and KAN to capture the dynamic progression of patient states.
@@
    Demonstrating TCKAN’s superior prediction accuracy and robustness over existing methods compared to existing methods, validated across two datasets.
@@ 2.2 Data Preprocessing
Based on the Sepsis-3 definition[33], we selected sepsis patients from the MIMIC-III and MIMIC-IV datasets who had a SOFA score of 2 or higher, as illustrated in Figure 2. Initially, we excluded patients under the age of 16 and those with corrupted data. We also eliminated patients with suspected infections occurring more than 24 hours before or after ICU admission, due to the ambiguous timing of their infection diagnoses, which could compromise the study’s accuracy. Furthermore, we excluded patients with ICU stays shorter than 24 hours, as their incomplete clinical data insufficiently support the model. Additionally, patients who underwent cardiac, thoracic, or vascular surgery were omitted because their unique surgical treatments and care requirements could adversely affect the model’s generalizability. For sepsis patients with multiple hospital admissions, we used data from their initial admission. We extracted a total of 10,567 patients from the MIMIC-III dataset (1,190 positive cases and 9,377 negative cases) and 11,540 patients from the MIMIC-IV dataset (1,274 positive cases and 10,266 negative cases).
@@
Using these criteria, we extracted patients’ constant and temporal data, along with diagnostic ICD coding information, from the MIMIC-III and MIMIC-IV databases. Constant data includes basic demographic information like gender, age, and weight, which provides essential background on the patients. Temporal data encompasses detailed physiological signs and laboratory test results recorded during the ICU stay, such as temperature, heart rate, blood pressure, respiratory rate, hemoglobin levels, and white blood cell count, reflecting the immediate health status and condition changes of the patients. Additionally, diagnostic ICD codes were transformed into vector representations via graph embedding methods, allowing for a comprehensive integration of patients’ constant characteristics, temporal features, and diagnostic information. The patient features extracted from the MIMIC-III and MIMIC-IV databases were obtained using SQL queries. Table 1 provides a detailed listing of the features that were selected and extracted using SQL methods from the temporal and constant data available in the MIMIC-III and MIMIC-IV databases.
@@
Temporal Feature Platelet, MCH, MCHC, Fibrinogen, Basophils, Lymphocytes, HCO3, Troponin-t, Sodium, Urine output, Bilirubin, HR, Lactic acid, Albumin, Calcium, Glucose, PH, Creatinine, Eosinophils, Total CO2, SBP, RBC, Chloride, ALT, Fraction inspired oxygen set, Alkaline phosphate, PaO2, GCS, PT, SpO2, Anion gap, DBP, Magnesium, BUN, WBC, Glucose, Temperature, MCV, Troponin-i, Potassium, Neutrophils, Hemoglobin, PTT, Lactate, PaCO2, FiO2, Hematocrit, AST, Monocytes, Phosphate, INR, MBP, RR
@@
In preprocessing the constant data extracted from patients, we employed specific strategies tailored to different types of missing data. For continuous variables, such as age and weight, missing values were imputed with zero to maintain data completeness and consistency without introducing bias, thereby ensuring that the model’s training is not negatively impacted. For categorical variables, such as gender and race, one-hot encoding was utilized, converting each category into a binary vector where each category is represented by a unique dimension. This method prevents the imposition of unintended ordinal relationships among categories, thereby avoiding any misleading effects on the model’s training.
For temporal data, we adhered to strict time window and interval strategies. Temporal data were extracted from the 24 hours preceding ICU admission and were sampled at 1-hour intervals. For each interval, if a single measurement was available, it was used directly; if multiple measurements were recorded, they were averaged to represent the feature level during that hour. In intervals where no measurements were recorded, values were marked as missing. After processing the constant and temporal features of the patients, we applied mean-variance normalization to the continuous variables within these features. We calculated the mean and standard deviation of each variable, using these parameters to adjust the data. This ensured that all continuous variables achieved a standardized form with zero mean and unit variance. Figure 3 illustrates the processing of temporal data. We will address the treatment of missing values in temporal data during subsequent model processing.
@@
$H^{(l+1)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}H^{(l)}W^{(l)}\right)$
@@ 3 Label
2.3 The Architecture of TCKAN
The TCKAN model architecture includes three components: (1) Temporal data processed by GRU-D (Gated Recurrent Unit with Decay) to generate hidden representations; (2) ICD and CCS codes analyzed via an attention mechanism to capture relationships and enrich diagnostic features; (3) Constant data processed through Kolmogorov–Arnold Networks (KAN). This network efficiently leverages constant feature information to extract high-level features via a multi-layer neural network. Subsequently, the hidden features from these three components are concatenated to form an integrated feature vector. This vector is further processed through a final KAN network to predict the patient’s sepsis mortality risk.
@@ 2.3 The Architecture of TCKAN
$D$ $=(D_{1},D_{2},\dots,D_{t},\dots,D_{T})\in\mathbb{R}^{T\times N}$
@@ 2.4 Implementation
In our research, we implemented and trained our model using the TensorFlow framework. During the training process, we systematically optimized several key parameters to improve the model’s performance including learning rate, batch size and number of iterations. We utilized the Adam optimizer to update model parameters and employed a learning rate decay strategy to dynamically adjust the learning rate, enhancing the model’s convergence stability. To address the significant imbalance between positive and negative samples, we applied oversampling techniques to maintain a consistent ratio of these samples across training, validation, and test sets. This method effectively mitigated model bias toward the majority class, thereby enhancing its generalization capabilities on imbalanced datasets. Moreover, we adopted five-fold cross-validation to validate the model’s stability and generalizability. By dividing the dataset into five mutually exclusive subsets, using one for validation and the remaining four for training—and repeating this process five times—we ensured the model’s consistent performance across various data splits. To prevent overfitting, we incorporated Dropout layers and monitored the validation set performance using early stopping techniques, halting training when no further reduction in validation loss was observed.
@@ 3.1 Performance Metrics
In this section, we describe the metrics employed to assess the TCKAN model’s performance. These metrics are vital for gauging the model’s accuracy in predicting mortality risk among ICU sepsis patients. The principal evaluation metrics include sensitivity, specificity, area under the receiver operating characteristic curve (AUROC), area under the precision-recall curve (AUPRC), and the Brier Score (BS). Each metric offers a unique perspective on the model’s performance, enabling a thorough analysis. Sensitivity, also known as recall, measures the proportion of positive samples correctly identified by the model. Specificity assesses the proportion of negative samples accurately recognized by the model. The AUROC evaluates the model’s overall classification ability across different thresholds; a value closer to 1 indicates stronger discriminative power. AUPRC is a metric that balances precision and recall, particularly useful in cases of imbalanced class distribution. The Brier Score gauges the accuracy of probabilistic predictions, with lower values signifying better accuracy. The specific formulas for these metrics are as follows:
@@ 3.2 Model Performance
We evaluated the TCKAN model’s performance on the MIMIC-III and MIMIC-IV datasets by comparing it with seven established baseline models: Xgboost[34], SVM, Random Forest (RF)[35], LGBM[17], LSTM[26], IseeU[36], and and Che et al.’s model GRU-D[30]. This comparison across various metrics such as sensitivity, specificity, AUROC, AUPRC, and Brier score allowed us to thoroughly assess the strengths and weaknesses of each model. Figure 6 illustrates the comparative performance, demonstrating that the TCKAN model generally outperforms the baseline models. Notably, on the MIMIC-IV dataset, the TCKAN model achieved AUROC and AUPRC values of 0.8807 and 0.5470, respectively, significantly surpassing its competitors. This indicates its superior accuracy and stability in distinguishing between positive and negative cases and handling imbalanced datasets. Additionally, the TCKAN model’s balanced performance in sensitivity (0.7875) and specificity (0.8153) underscores its effectiveness in identifying patients at different levels of risk. A lower Brier score of 0.0696 reflects the model’s precision in probability predictions. Figure 7, showing the ROC curves, further highlights the TCKAN model’s enhanced ability to differentiate between outcomes, with its curve closer to the top-left corner, symbolizing higher prediction accuracy across various thresholds.
@@
Figure 7: Receiver Operating Characteristic (ROC) curves, employing 5-fold cross-validation on MIMIC-III and MIMIC-IV datasets, display the performance of various models. These include GRUD, ISEEU, LightGBM, Logistic Regression (LR), Long Short-Term Memory (LSTM), Random Forest (RF), Support Vector Machine (SVM), TCKAN, and XGBoost.
To further evaluate the robustness of the TCKAN model, we conducted parameter sensitivity experiments focusing on the learning rate (LR) and batch size, two crucial hyperparameters that influence model performance. We systematically varied these parameters, setting the learning rate at 0.0001, 0.001, 0.01, and 0.1, and the batch size at 16, 32, 64, and 128, to observe their effects on model performance under various combinations. Figure 8 illustrates these performance changes through a three-dimensional graph that plots AUROC values against variations in learning rate and batch size. The results indicate that the TCKAN model achieves optimal performance with a learning rate of approximately 0.001 and a batch size of 64. Notably, excessively high or low learning rates can degrade performance, while very small batch sizes can introduce instability during training, and overly large batch sizes may reduce training speed and impair generalization.
Figure 8: The impact of varying batch sizes and learning rates on the performance of the TCKAN model across MIMIC-III and MIMIC-IV datasets. Each bar represents the Area Under the Receiver Operating Characteristic Curve (AUROC), with colors indicating different AUROC values as batch size and learning rate change.